# MINDS: Detection of an inner gas disk caused by evaporating bodies around HD 172555

### v2

@@ 3.3 Origin of atomic gas lines
In HD 172555 the lines appear to be a direct tracer of evaporating bodies or dust releasing material into the gas phase. The analysis of the line broadening (see Table 2) shows that the lines originate very close to the star. The equilibrium temperature of an irradiated body without heat redistribution is approximated by
where , is the albedo, the separation from the star, and the Stefan-Boltzman constant. Most solar system asteroids and comets tend to have low albedos (0.1, e.g., Wright et al., 2016). This means that a temperature of about 720 K, enough to break apart troilite (FeS, Lauretta et al., 1995), which is the simplest refractory compound that can release sulfur, is reached at a separation of around 0.8 au (assuming ). At 0.3 au temperatures approach 1200 K, which is high enough to break apart most silicates (Lenzuni et al., 1995). It therefore stands to reason that the gas is probing the region in which rocky material can sublimate.
@@ 3.3.1 Impact of stellar radiation on the stability of the gas reservoir
Once gas is released, it is important to assess whether the detected species can form a stable reservoir in the presence of the strong stellar radiation field of the A-type star. Atoms and ions, can experience acceleration by absorbing stellar photons, which may remove them from the inner disk.
Chlorine does not experience strong radiation pressure from the star, as neither its neutral nor ionized form posses abundant and strong transitions in the UV or optical range, where the photon flux of the A-type host star is high (Lehtmets et al., 2024). Fernández et al. (2006) showed that while ionized species of iron and nickel do experience outward acceleration due to radiation pressure in Pic, but could be self-breaking due to enhanced carbon abundance. The newly discovered neutral sulfur may also be affected by radiative blowout, as it has a in Lehtmets et al. for an A6V star, and in Fernández et al. (2006) for Pic. We expect similar chromospheric activity in HD 172555 compared to Pic (Grady et al., 2018), and therefore a radiation field of comparable strength (also see Appendix H for an analysis of the XUV and X-ray fields). Therefore, an additional breaking mechanism, like the self-breaking due to carbon, may also be necessary to explain the strong and stable sulfur detection (see Section 3.3.3). Circumstellar carbon has been observed in absorption around HD 172555 by Grady et al. (2018).
@@ 3.3.3 Neutral sulfur S i
This is the first time that atomic sulfur is reported in the mid-IR spectrum of a debris disk. The line is located at 25.249 m and is the strongest emission line in the spectrum. It is the only line that appears to be asymmetric in the amplitude of the two Gaussian components (see zoom-in in Fig. 1 and the 2- residuals in the symmetric Keplerian model fit Fig. 8 e), potentially hinting at a stronger blue shifted component.
From the archival Spitzer/IRS LH () we succeeded to detect the [S i] line (see Appendix A). None of the other lines are detectable in Spitzer due to insufficient signal-to-noise ratio and resolution. This detection of the line in JWST after 20-years, at a similar or slightly higher flux level, means that not only is the dust emission stable on these time-scales (see Sect. 3.6 and Su et al., 2020), but also the sulfur gas emission. The slight increase in the silicate emission (Fig. 1) may also be consistent with a slight increase in sulfur emission of about 20% (see Appendix A), but this is hard to conclusively show given the larger uncertainties on the Spitzer data.
@@
In the solar system the organic sulfur molecular content has been measured directly in samples of the asteroid 162173 Ryugu, where the total sulfur abundance was between 2.5–6.2 weight percent (Yoshimura et al., 2023). Many carbonaceous chondrites found on Earth have similar bulk abundances between 3–6 weight percent (Alexander et al., 2022).
The near-Sun asteroid (3200) Phaeton (Jewitt & Li, 2010), the likely parent body of the Geminid meteoroid stream, could be a good analog to study evaporating exo-asteroids. With a perihelion distance of 0.14 au, it undergoes periodic heating episodes up to 750 K, showing comet-like activity. A recent laboratory study by Suttle et al. (2024) showed that the retention of sulfur-bearing gas may play a crucial role in limiting thermal decomposition, allowing FeS to survive repeated heating cycles. This could provide a viable mechanism for a population of bodies to survive repeated non-catastrophic thermal decomposition events and release volatiles over an extended period of time.
@@ 3.3.4 Neutral chlorine Cl i
Atomic chlorine is seen in the mid-IR spectrum of a debris disk for the first time. The line, located at 11.333 m is the second strongest emission line in the spectrum after neutral sulfur at 25.249 m. The [Cl i] line has recently been found to trace jets and shocked knots in protostars (Tychoniec et al., 2024; Nisini et al., 2024; Caratti o Garatti et al., 2024). In the solar system, evidence for chlorine has been found in ammonium salts (NH4Cl, “salmiac salt”) in comet 67P (Altwegg et al., 2020), which is thought to sublimate at a temperature of around 473 K (Clementi & Gayles, 1967). In the coma of comet 67P the volatile fraction of chlorine relative to oxygen is (Dhooghe et al., 2017), whereas in chondrites the ratio is (Lodders, 2010). Chlorine bearing salts have also been directly measured in samples of asteroid 162173 Ryugu (Yoshimura et al., 2023). The chlorine-to-oxygen ratio could therefore be a potential indicator for the type of objects that produce the observed chlorine signature. Neutral oxygen has been detected in emission with Herschel around HD 172555 (see Table 1). However, it is unclear what radial separation the emission traces. If a giant impact is the origin of the oxygen, then the probed oxygen reservoir may not be co-located with the measured chlorine. However, the oxygen could also be released over time from evaporating rocky material similarly to the observed chlorine.
@@ 3.4 Physical conditions of the Fe ii gas
The detection of multiple transitions of [Fe ii] provides unique access to the local conditions in the reservoir of atomic gas. Because of the high critical density and the expected high electron abundance, the line ratio is sensitive to the local electron density. Following the approach outlined in Caratti o Garatti et al. (2024), we derive an estimate for the electron density () and temperature () based on the observed [Fe ii] line ratios and upper limits. The model is based on a non-local thermal equilibrium (nLTE) excitation model presented in Giannini et al. (2013) and updated for MIRI-MRS transitions at low excitation temperatures. It assumes collisional excitation/de-excitation with electrons and spontaneous radiative decay. This is shown in Fig. 2, including, for comparison, the line fluxes of Pic from Wu et al. (2025), namely and erg s-1 cm2 for the 17.9 and 26 m lines respectively. For [Fe ii] at 5.3 m we can only derive an upper limit of erg s-1 cm2 (corresponding to 1-). The values for HD 172555 are listed in Table 1. For [Fe ii] at 5.3 m we determine a flux of erg s-1 cm2, following the methodology outlined in Appendix B.2.
The similar flux of the Fe ii lines at 26 m and 17.9 m indicates a high electron temperature , whereas the low flux of [Fe ii] at 5.3 m only provides a rough constraint of the electron density to cm-3. Pic has a K and a lower limit of cm-3. In general HD 172555’s gas has higher electron temperature and density than the densest region (B3) of the atomic jet extending from the protostar HH 211: cm-3 and K (Caratti o Garatti et al., 2024). All of this is consistent with the overall picture of a close-in gas disk caused by evaporating dust/asteroids/exocomets, as opposed to the shock excitation in an atomic jet environment. The estimated absolute values may not reflect the real temperatures and densities due to the neglected UV excitation which should play a large role due to the proximity to the A-type star. It is also uncertain how much of the UV radiation is shielded by dust and gas close to the star. However, qualitatively the relative comparison of HD 172555 to Pic is still informative.
@@ 3.5 Molecular lines
It has been suggested by Lisse et al. (2009) that SiO gas could explain a part of the mid-IR excess feature that was hard to explain with their dust model. With the spectral resolution of JWST MIRI/MRS, we would have been able to resolve SiO lines if the gas were present in large enough quantity to explain the quasi-continuum feature invoked by Lisse et al. (2009). However, we do not see any clear indication of SiO gas emission lines in the range from 7–9 m. This is consistent with a previous non-detection of SiO gas using the APEX submillimeter telescope (Wilson et al., 2016).
Either SiO is not released in sufficient quantities in the gas phase, or as Johnson et al. (2012) have argued, could be photo-dissociated. Our observations provide further indication that SiO is not present in the gas phase in significant amounts.
@@ 3.6 Dust
The dust properties have been investigated previously based on the Spitzer data (e.g. Lisse et al., 2009; Johnson et al., 2012; Morlok et al., 2012). We find that our MRS spectra are broadly consistent with the Spitzer measurement. Starting from around 9 m the silicate feature in our data is about 5% brighter than previously measured in 2006 and 2008 (see Fig. 1). This is still consistent with the previous measurements within the calibration uncertainties of the Spitzer data. However, the offsets are primarily seen in the silicate feature and not the stellar photosphere, indicating that the change may be real. Other monitoring campaigns at 3.6 and 4.5 m (Su et al., 2020) showed % variability over the last 20 years. IRAS observations at 12 and 25 m from 1983 are compatible with WISE observations at 12 and 22 m in 2010 within 4% (Johnson et al., 2012). This means that the flux seems to be fairly stable over the baseline of 40 years.
@@ 4 Discussion
The new MIRI/MRS data adds a new fascinating layer to our understanding of the system’s architecture. Previously, ALMA has seen the warm (169 K) dust continuum extending out to 10 au, similar in spatial distribution to the CO gas (Schneiderman et al., 2021). The bulk gas and dust detected by ALMA are expected to be at separation between around 4–7 au, which is consistent with the location of the warm dust as determined from the silicate feature (Lisse et al., 2009; Johnson et al., 2012). Small grains of sizes from 1.25–3.9 m are also detected out to about 10 au in resolved polarimetric high-contrast observation (Engler et al., 2018) and modeling based on unresolved polarimetry (Marshall et al., 2020). Our newly detected gas disk with JWST MIRI is much closer-in than the bulk of the gas and dust previously observed. Ertel et al. (2014) have reported a tentative detection of hot dust around HD 172555 with VLTI/PIONIER, which, if real, may be produced by the same mechanism as the gas reported in our study.
@@ 4.1 Is the atomic gas tracing a different population of bodies than in Pictoris?
We have reduced and continuum‐subtracted the MIRI/MRS data of Pic (Worthen et al., 2024) using the same methodology as for HD 172555. The continuum‐subtracted spectra, scaled to the distance of HD 172555, are shown in the Appendix E (Fig. 5). Although both stars belong to the same moving group with similar ages and spectral types, they have different system architectures and debris disk properties. The differences in detected emission lines indicate that we may be probing distinct populations of bodies.
@@ 4.1.1 Line Profiles and Physical Conditions
Wu et al. (2025) report the detection of [Fe ii] lines at 17.9 and 26 m in Pictoris, which we also observe in HD 172555. Our analysis shows that these lines in Pic (see Table 5) also appear Doppler broadened, but show velocities roughly half those in HD 172555. This implies that the bulk of the [Fe ii] emission originates at larger separations ( au; see Appendix Fig. 13), consistent with Wu et al. (2025) assessment that this gas may be part of the stable gas disk component. The conclusion that the gas probes larger separations is reinforced by the [Fe ii] line ratios (Fig. 2; Section 3.4), which indicate that the electron temperature in HD 172555 is about twice as high as in Pic.
@@ 4.1.2 Implications for the Nature of the Evaporating Bodies
(i) Perturbed asteroid-like bodies: Bodies are scattered into close stellar encounters, crossing different sublimation thresholds that trigger gas release. This mechanism is supported by the detection of falling evaporating bodies (FEBs) (Kiefer et al., 2014; Grady et al., 2018), with Grady et al. (2018) noting that the simultaneous detection of carbon and Ca ii ions in absorption suggests a lower volatile-to-refractory ratio compared to solar system comets. However, the observed stability of the inner gas disk in HD 172555 indicates that continuous or quasi-continuous gas production from stable close-in material may be required, rather than solely transient events (see Section 4.1.3).
(ii) In-situ collisional cascades: Collisions of bodies near the star continuously produce small dust that is rapidly evaporated by stellar heating. However, maintaining a significant population of parent bodies at such close separations is challenging due to their short collisional time scale (on the order of thousands of years), compared to the system age 20 Myr (Wyatt et al., 2007; Kral et al., 2017; Wu et al., 2025).
(iii) Outgassing of a large body: A single large object, for instance an evaporating rocky world such as BD+054868 Ab (Hon et al., 2025), may continuously supply gas through outgassing and thermal decomposition.
(iv) Evaporation of dust: Su et al. (2020) suggested that the giant impact invoked to explain the systems’ mineralogy and CO content could generate an optically thick clump of mm-sized vapor condensates. This clump would provide shielding from radiation pressure blowout, while sub-m-sized grains are produced in large quantities in the clump. They would persist after the clump disperses into an optically thin ring due to being small
enough that their optical properties make them resistant to blow-out. Poynting–Robertson drag would then slowly cause them to approach the star on time scales of 1.6–3.6 yr, where they sublimates and releases gas.
In Pic the known gas components form multiple distinct reservoirs. Transient absorption features in UV and optical spectra, attributed to individual falling evaporating bodies, indicate separations at transit of 10–20 R∗ (0.05–0.1 au) (Kiefer et al., 2014). Originally, these bodies were thought to originate from a reservoir located around 4–5 au from the star. However, a new dynamical analysis of the system by Beust et al. (2024), accounts for the more recently discovered planet Pic c at 2.7 au and concludes that only a population of bodies within 1.5 au would remain stable enough to supply FEBs long-term. This is broadly consistent with TESS transit observations of small objects in the system (Heller, 2024).
@@
In the MIRI MRS data of Pic we locate the bulk of the iron and nickle gas at separations beyond 1 au, suggesting it is more likely associated with a persistent gas disk component rather than the dynamically short-lived gas released by FEBs seen in absorption. The orbital velocity of FEBs near periastron are high, making it challenging for FEB-produced gas alone to form a persistent inner disk. This distinction underscores the potentially different mechanisms maintaining gas in the two systems: in HD 172555, an ongoing supply from rocky, very close-in evaporating bodies and/or dust, whereas in Pic, a mixture of short-lived exocometary gas and an extended, stable, carbon-rich disk.
@@
Since argon is highly volatile and not easily incorporated into solids, its presence in Pic indicates that the bodies releasing it formed in cold regions (T 35 K) where argon could be trapped in ices (e.g. Bar-Nun et al., 1988) or clathrate hydrates (e.g. Lunine & Stevenson, 1985). This is consistent with Pic hosting an extended, cold debris disk and exhibiting disk substructures — such as the Cat’s Tail and a misaligned sub-disk—that may arise from collisions in the outer system (Rebollido et al., 2024).
The pronounced neutral chlorine and sulfur emission features in HD 172555, which are not detected in Pic, may arise from differences in ionization conditions or production mechanisms. In Pic, a higher UV/X-ray flux or a difference in the disk structure could lead to lower abundances of neutral Cl and S. We find an 50% higher FUV flux in archival Pic HST/COS spectra compared to HD 172555. The XUV and X-ray fields of both sources could be similar based on their spectral types and ages; however, HD 172555 is not clearly detected in existing X-ray archival data (see Appendix H). The atomic species observed in the MRS spectra have widely varying ionization potentials (Appendix Table 9). The observed pattern of detections and non-detections in the two systems is not straightforwardly explained by a simple change in the number of hardness of ionizing photons.

### v3

@@ 3.3 Origin of atomic gas lines
In HD 172555, the lines appear to be a direct tracer of evaporating bodies or dust releasing material into the gas phase. The analysis of the line broadening (see Table 2) shows that the lines originate very close to the star. The equilibrium temperature of an irradiated body without heat redistribution is approximated by
where , is the albedo, is the separation from the star, and is the Stefan–Boltzman constant. Most Solar System asteroids and comets tend to have low albedos (0.1, e.g., Wright et al., 2016). This means that a temperature of about 720 K, enough to break apart troilite (FeS; Lauretta et al., 1995), which is the simplest refractory compound that can release sulfur, is reached at a separation of around 0.8 au (assuming ). At 0.3 au, temperatures approach 1200 K, which is high enough to break apart most silicates (Lenzuni et al., 1995). It therefore stands to reason that the gas is probing the region in which rocky material can sublimate.
@@ 3.3.1 Impact of stellar radiation on the stability of the gas reservoir
Once gas is released, it is important to assess whether the detected species can form a stable reservoir in the presence of the strong stellar radiation field of the A-type star. Atoms and ions can experience acceleration by absorbing stellar photons, which may remove them from the inner disk.
Chlorine does not experience strong radiation pressure from the star, as neither its neutral nor ionized form posses abundant and strong transitions in the UV or optical range, where the photon flux of the A-type host star is high (Lehtmets et al., 2024). Fernández et al. (2006) showed that while ionized species of iron and nickel do experience outward acceleration due to radiation pressure in Pic, they could be self-breaking due to enhanced carbon abundance. The newly discovered neutral sulfur may also be affected by radiative blowout, as it has a in Lehtmets et al. 2025, in preparation, for an A6V star, and in Fernández et al. (2006) for Pic. We expect similar chromospheric activity in HD 172555 compared to Pic (Grady et al., 2018), and therefore a radiation field of comparable strength (also see Appendix H for an analysis of the extreme-UV (XUV) and X-ray fields). Therefore, an additional breaking mechanism, like the self-breaking due to carbon, may also be necessary to explain the strong and stable sulfur detection (see Section 3.3.3). Circumstellar carbon has been observed in absorption around HD 172555 by Grady et al. (2018).
@@ 3.3.3 Neutral sulfur S i
This is the first time that atomic sulfur is reported in the mid-IR spectrum of a debris disk. The line is located at 25.249 m and is the strongest emission line in the spectrum. It is the only line that appears to be asymmetric in the amplitude of the two Gaussian components (see zoom-in in Fig. 1 and the 2- residuals in the symmetric Keplerian model fit in the Appendix Fig. 8 e), potentially hinting at a stronger blue shifted component.
From the archival Spitzer/IRS LH () we succeeded to detect the [S i] line (see Appendix A). None of the other lines are detectable in Spitzer due to insufficient signal-to-noise ratio and resolution. This detection of the line in JWST after 20-years, at a similar or slightly higher flux level, means that not only is the dust emission stable on these timescales (see Sect. 3.6 and also Su et al., 2020), but so too is the sulfur gas emission. The slight increase in the silicate emission (Fig. 1) may also be consistent with a slight increase in sulfur emission of about 20% (see Appendix A), but this is hard to conclusively show given the larger uncertainties on the Spitzer data.
@@
In the Solar System, the organic sulfur molecular content has been measured directly in samples of the asteroid 162173 Ryugu, where the total sulfur abundance was between 2.5 and 6.2 weight percent (Yoshimura et al., 2023). Many carbonaceous chondrites found on Earth have similar bulk abundances between 3 and 6 weight percent (Alexander et al., 2022).
The near-Sun asteroid (3200) Phaeton (Jewitt & Li, 2010), the likely parent body of the Geminid meteoroid stream, could be a good analog to study evaporating exo-asteroids. With a perihelion distance of 0.14 au, it undergoes periodic heating episodes up to 750 K, showing comet-like activity. A recent laboratory study by Suttle et al. (2024) showed that the retention of sulfur-bearing gas may play a crucial role in limiting thermal decomposition, allowing FeS to survive repeated heating cycles. This could provide a viable mechanism for a population of bodies to survive repeated noncatastrophic thermal decomposition events and release volatiles over an extended period of time.
@@ 3.3.4 Neutral chlorine Cl i
Atomic chlorine is seen in the mid-IR spectrum of a debris disk for the first time. The line, located at 11.333 m is the second strongest emission line in the spectrum after neutral sulfur at 25.249 m. The [Cl i] line has recently been found to trace jets and shocked knots in protostars (Tychoniec et al., 2024; Nisini et al., 2024; Caratti o Garatti et al., 2024). In the Solar System, evidence for chlorine has been found in ammonium salts (NH4Cl, “salmiac salt”) in comet 67P (Altwegg et al., 2020), which is thought to sublimate at a temperature of around 473 K (Clementi & Gayles, 1967). In the coma of comet 67P, the volatile fraction of chlorine relative to oxygen is (Dhooghe et al., 2017), whereas in chondrites the ratio is (Lodders, 2010). Chlorine bearing salts have also been directly measured in samples of asteroid 162173 Ryugu (Yoshimura et al., 2023). The chlorine-to-oxygen ratio could therefore be a potential indicator for the type of objects that produce the observed chlorine signature. Neutral oxygen has been detected in emission with Herschel around HD 172555 (see Table 1). However, it is unclear what radial separation the emission traces. If a giant impact is the origin of the oxygen, then the probed oxygen reservoir may not be colocated with the measured chlorine. However, the oxygen could also be released over time from evaporating rocky material similarly to the observed chlorine.
@@ 3.4 Physical conditions of the Fe ii gas
The detection of multiple transitions of [Fe ii] provides unique access to the local conditions in the reservoir of atomic gas. Because of the high critical density and the expected high electron abundance, the line ratio is sensitive to the local electron density. Following the approach outlined in Caratti o Garatti et al. (2024), we derive an estimate for the electron density () and temperature () based on the observed [Fe ii] line ratios and upper limits. The model is based on a nonlocal thermal equilibrium (nLTE) excitation model presented in Giannini et al. (2013) and updated for MIRI MRS transitions at low excitation temperatures. It assumes collisional excitation/de-excitation with electrons and spontaneous radiative decay. This is shown in Fig. 2, including, for comparison, the line fluxes of Pic from Wu et al. (2025), namely and erg s-1 cm2 for the 17.9 and 26 m lines, respectively. For [Fe ii] at 5.3 m we can only derive an upper limit of erg s-1 cm2 (corresponding to 1-). The values for HD 172555 are listed in Table 1. For [Fe ii] at 5.3 m we determine a flux of erg s-1 cm2, following the methodology outlined in Appendix B.2.
The similar flux of the Fe ii lines at 26 m and 17.9 m indicates a high electron temperature , whereas the low flux of [Fe ii] at 5.3 m only provides a rough constraint of the electron density to cm-3. Pic has a K and a lower limit of cm-3. In general, HD 172555’s gas has higher electron temperature and density than the densest region (B3) of the atomic jet extending from the protostar HH 211: cm-3 and K (Caratti o Garatti et al., 2024). All of this is consistent with the overall picture of a close-in gas disk caused by evaporating dust/asteroids/exocomets, as opposed to the shock excitation in an atomic jet environment. The estimated absolute values may not reflect the real temperatures and densities due to the neglected UV excitation, which should play a large role due to the proximity to the A-type star. It is also uncertain how much of the UV radiation is shielded by dust and gas close to the star. However, qualitatively, the relative comparison of HD 172555 to Pic is still informative.
@@ 3.5 Molecular lines
It has been suggested by Lisse et al. (2009) that SiO gas could explain a part of the mid-IR excess feature that was hard to explain with their dust model. With the spectral resolution of JWST MIRI MRS, we would have been able to resolve SiO lines if the gas were present in large enough quantities to explain the quasi-continuum feature invoked by Lisse et al. (2009). However, we do not see any clear indication of SiO gas emission lines in the range from 7–9 m. This is consistent with a previous nondetection of SiO gas using the APEX submillimeter telescope (Wilson et al., 2016).
Either SiO is not released in sufficient quantities in the gas phase, or as Johnson et al. (2012) have argued, it could be photodissociated. Our observations provide further indication that SiO is not present in the gas phase in significant amounts.
@@ 3.6 Dust
The dust properties have been investigated previously based on the Spitzer data (e.g., Lisse et al., 2009; Johnson et al., 2012; Morlok et al., 2012). We find that our MRS spectra are broadly consistent with the Spitzer measurement. Starting from around 9 m the silicate feature in our data is about 5% brighter than previously measured in 2006 and 2008 (see Fig. 1). This is still consistent with the previous measurements within the calibration uncertainties of the Spitzer data. However, the offsets are primarily seen in the silicate feature and not the stellar photosphere, indicating that the change may be real. Other monitoring campaigns at 3.6 and 4.5 m (Su et al., 2020) showed 0.5% variability over the last 20 years. IRAS observations at 12 and 25 m from 1983 are compatible with WISE observations at 12 and 22 m in 2010 within 4% (Johnson et al., 2012). This means that the flux seems to be fairly stable over the baseline of 40 years.
@@ 4 Discussion
The new MIRI MRS data adds a new fascinating layer to our understanding of the system’s architecture. Previously, ALMA has seen the warm (169 K) dust continuum extending out to 10 au, similar in spatial distribution to the CO gas (Schneiderman et al., 2021). The bulk gas and dust detected by ALMA are expected to be at separation between around 4 and 7 au, which is consistent with the location of the warm dust as determined from the silicate feature (Lisse et al., 2009; Johnson et al., 2012). Small grains of sizes from 1.25–3.9 m are also detected out to about 10 au in resolved polarimetric high-contrast observation (Engler et al., 2018) and modeling based on unresolved polarimetry (Marshall et al., 2020). Our newly detected gas disk with JWST MIRI is much closer-in than the bulk of the gas and dust previously observed. Ertel et al. (2014) reported a tentative detection of hot dust around HD 172555 with VLTI/PIONIER, which, if real, may be produced by the same mechanism as the gas reported in our study.
@@ 4.1 Is the atomic gas tracing a different population of bodies than in Pictoris?
We have reduced and continuum‐subtracted the MIRI MRS data of Pic (Worthen et al., 2024) using the same methodology as for HD 172555. The continuum‐subtracted spectra, scaled to the distance of HD 172555, are shown in the Appendix E (Fig. 5). Although both stars belong to the same moving group with similar ages and spectral types, they have different system architectures and debris disk properties. The differences in detected emission lines indicate that we may be probing distinct populations of bodies.
@@ 4.1.1 Line Profiles and Physical Conditions
Wu et al. (2025) reported the detection of [Fe ii] lines at 17.9 and 26 m in Pictoris, which we also observe in HD 172555. Our analysis shows that these lines in Pic (see Table 5) also appear Doppler broadened, but show velocities roughly half of those in HD 172555. This implies that the bulk of the [Fe ii] emission originates at larger separations (0.8 au; see Appendix Fig. 13), consistent with the assessment of Wu et al. (2025) that this gas may be part of the stable gas disk component. The conclusion that the gas probes larger separations is reinforced by the [Fe ii] line ratios (Fig. 2; Section 3.4), which indicate that the electron temperature in HD 172555 is about twice as high as in Pic.
@@ 4.1.2 Implications for the Nature of the Evaporating Bodies
(i) Perturbed asteroid-like bodies. Bodies are scattered into close stellar encounters, crossing different sublimation thresholds that trigger gas release. This mechanism is supported by the detection of falling evaporating bodies (FEBs) (Kiefer et al., 2014; Grady et al., 2018), with Grady et al. (2018) noting that the simultaneous detection of carbon and Ca ii ions in absorption suggests a lower volatile-to-refractory ratio compared to Solar System comets. However, the observed stability of the inner gas disk in HD 172555 indicates that continuous or quasi-continuous gas production from stable close-in material may be required, rather than solely transient events (see Section 4.1.3).
(ii) In situ collisional cascades. Collisions of bodies near the star continuously produce small dust that is rapidly evaporated by stellar heating. However, maintaining a significant population of parent bodies at such close separations is challenging due to their short collisional time scale (on the order of thousands of years), compared to the system age 20 Myr (Wyatt et al., 2007; Kral et al., 2017; Wu et al., 2025).
(iii) Outgassing of a large body. A single large object, for instance an evaporating rocky world such as BD+054868 Ab (Hon et al., 2025), may continuously supply gas through outgassing and thermal decomposition.
(iv) Evaporation of dust. Su et al. (2020) suggested that the giant impact invoked to explain the systems’ mineralogy and CO content could generate an optically thick clump of mm-sized vapor condensates. This clump would provide shielding from radiation pressure blowout, while submicron-sized grains are produced in large quantities in the clump, which can explain the mid-IR spectral feature. They would persist after the clump disperses into an optically thin ring due to being small
enough that their optical properties make them resistant to blowout. On the other hand, Poynting–Robertson drag will slowly cause the millimeter-sized grains to approach the star on time scales of 1.6–3.6 yr, where they would sublimates and releases gas.
In Pic the known gas components form multiple distinct reservoirs. Transient absorption features in UV and optical spectra, attributed to individual FEBs, indicate separations at transit of 10–20 R∗ (0.05–0.1 au) (Kiefer et al., 2014). Originally, these bodies were thought to originate from a reservoir located around 4–5 au from the star. However, a new dynamical analysis of the system by Beust et al. (2024), has accounted for the more recently discovered planet Pic c at 2.7 au and concluded that only a population of bodies within 1.5 au would remain stable enough to supply FEBs long-term. This is broadly consistent with TESS transit observations of small objects in the system (Heller, 2024).
@@
In the MIRI MRS data of Pic we locate the bulk of the iron and nickle gas at separations beyond 1 au, suggesting it is more likely associated with a persistent gas disk component rather than the dynamically short-lived gas released by FEBs seen in absorption. The orbital velocity of FEBs near periastron is high, making it challenging for FEB-produced gas alone to form a persistent inner disk. This distinction underscores the potentially different mechanisms maintaining gas in the two systems: in HD 172555, an ongoing supply from rocky, very close-in evaporating bodies and/or dust, whereas in Pic, a mixture of short-lived exocometary gas and an extended, stable, carbon-rich disk.
@@
Since argon is highly volatile and not easily incorporated into solids, its presence in Pic indicates that the bodies releasing it formed in cold regions (T 35 K) where argon could be trapped in ices (e.g., Bar-Nun et al., 1988) or clathrate hydrates (e.g., Lunine & Stevenson, 1985). This is consistent with Pic hosting an extended, cold debris disk and exhibiting disk substructures—such as the Cat’s Tail and a misaligned subdisk—that may arise from collisions in the outer system (Rebollido et al., 2024).
The pronounced neutral chlorine and sulfur emission features in HD 172555, which are not detected in Pic, may arise from differences in ionization conditions or production mechanisms. In Pic, a higher UV/X-ray flux or a difference in the disk structure could lead to lower abundances of neutral Cl and S. We find an 50% higher FUV flux in archival Pic HST/COS spectra compared to HD 172555. The XUV and X-ray fields of both sources could be similar based on their spectral types and ages; however, HD 172555 is not clearly detected in existing X-ray archival data (see Appendix H). The atomic species observed in the MRS spectra have widely varying ionization potentials (Appendix Table 9). The observed pattern of detections and nondetections in the two systems is not straightforwardly explained by a simple change in the number of hardness of ionizing photons.